OpenStudy (anonymous):

Write the equation for a circle with center (-5,3) and radius...

OpenStudy (anonymous):

Radius would be \[\sqrt{3}\]

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf (x-h)^2+(y-k)^2=r^2 }\) h is the x coordinate of the center k is the y coordinate of the center (r^2 is radius to the second power.)

OpenStudy (anonymous):

\[(x--5)^{2}+(y+3)^{2}=?\]

OpenStudy (anonymous):

@SolomonZelman I don't know what the last part would be

OpenStudy (anonymous):

oh.. omg im dumb im sorry nevermind

OpenStudy (solomonzelman):

I am going to give you a hint \(\Large\color{purple}{ \bf \sqrt{3} ^{~~2} =3}\)

OpenStudy (solomonzelman):

Okay, can you write the equation of the circle, now ?

OpenStudy (anonymous):

Yeah, I got it. Thank you so much for all of your help!

OpenStudy (solomonzelman):

Can you write in here, to make sure you are correct? I would prefer you to do that, but if you don't want to, it's fine either way.